OpenStudy (anonymous):

solve each equation 36 divided by =-6

myininaya (myininaya):

divided by ? 36/?=-6 36=-6? 36/(-6)=? -6=?

OpenStudy (anonymous):

itss \[36\div x=-6\]

OpenStudy (anonymous):

x=-6\36 x=-1/6

OpenStudy (anonymous):

But i have to write in in a equation

OpenStudy (anonymous):

so we have 36/x = -6 multiply each side by x to get -6x=36 divide each side by -6 to get x=36/-6 and x=-6 usually that is all they are asking for

myininaya (myininaya):

x=-6 is an equation

OpenStudy (anonymous):

indeed -- but it is also the solution to the equation 36/x=-6 I guess you could look at it a simplifying the main equation...I'm not sure what you are asking there.

OpenStudy (anonymous):

your suposed to figure out what x it to make -6\

myininaya (myininaya):

we did x=-6

OpenStudy (anonymous):

if you are really against the idea of the answer being an equation, you could go all commando on the answer and just put -6 and circle it.

myininaya (myininaya):

lol

OpenStudy (anonymous):

i wasnt understanding it sorry lol

myininaya (myininaya):

x=-6 is just saying if x is -6 then thats when the left hand side is the same as the right hand side

myininaya (myininaya):

36/-6=-6 is true right?

OpenStudy (anonymous):

now i got -8x=56

myininaya (myininaya):

you want to solve for x divide both sides by -8 x=56/(-8)=-7 is -8(-7) the same as 56

myininaya (myininaya):

say yes lol

OpenStudy (anonymous):

yeayes lol

myininaya (myininaya):

so the answer is -7 or you can be formal about it and say x=-7 i have to get ready for bed good luck

OpenStudy (anonymous):

?=-6 in every case. 36 divide by -6 is -6. The product of 6 and -6 is -36. The third item is the same as the first. -6=-6

OpenStudy (anonymous):

The product of 6 and -6 is -36. should have been written as follows: The product of -6 and -6 is 36.